OpenStudy (anonymous):

one number is 2 times a first number. a third number is 100 more than the first number. if the sum of the three numbers is 416, find the numbers?

OpenStudy (anonymous):

2n+100=416 I believe that is you're equation

OpenStudy (acxbox22):

actually the first number is n, the second 2n, and the third n+100 so when you add the three numbers you get the equation n+2n+n+100=416

OpenStudy (anonymous):

first number = n second number = 2n third number = n + 100 (n) + (2n) + (n + 100) = 416 --- combine like terms 4n + 100 = 416 --- subtract 100 from both sides 4n = 416 - 100 4n = 316 --- divide both sides by 4 n = 316/4 n = 79 2n = 2(79) = 158 n + 100 = 79 + 100 = 179 your numbers are : 79,158, and 179